### v3

# A bound for internal radii of stable manifolds in terms of Lyapunov exponents

Abstract.
We find some bounds for the internal radii of stable and unstable manifolds of points in terms of their Lyapunov exponents under the assumption of the existence of a dominated splitting.
The author has been supported by the NSFC11871262, NSFC11871394, and NSFC12250710130 funds.
1. Introduction
Let $M$ be a closed connected Riemannian manifold. Let $f\in\operatorname{Diff}^{1+}(M)$. For any given invertible operator AA, let us define its conorm $m(A)$ by $\|A^{-1}\|^{-1}$. We say that $f$ admits a $\gamma$-dominated splitting with $\gamma>0$ if there is a DfDf-invariant splitting $TM=E^{-}\oplus E^{+}$ such that:
(1.1) $\|Df_{-}(x)\|<e^{-2\gamma}m(Df_{+}(x)),$
where $Df_{\pm}(x)=Df(x)|_{E^{\pm}}$.
The diffeomorphism $f$ has a dominated splitting if it admits a $\gamma$-dominated splitting for some $\gamma>0$.
We define the following exponents:
$LE_{-}(x)=\limsup_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\log\|Df_{-}(f^{k}(x))\|$
and
$LE_{+}(x)=\liminf_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\log m(Df_{+}(f^{k}(x)))$
We define the Pesin stable and unstable manifolds by:
$W^{-}(x)=\{y\in M:\limsup_{n\to\infty}\frac{1}{n}\log d(f^{n}(x),f^{n}(y))<0\}$
and
$W^{+}(x)=\{y\in M:\limsup_{n\to\infty}\frac{1}{n}\log d(f^{-n}(x),f^{-n}(y))<0\}$
For any $x\in M$, we define the maximal internal radius of $W^{\pm}(x)$ by
$R_{\pm}(x)=\inf\{\operatorname{length}(\alpha):\alpha(0)=x,\,\alpha(1)\in\partial W^{\pm}(x),\,\alpha(t)\in W^{\pm}(x)\,\forall t\in[0,1)\}.$
A function $\phi:M\to\mathbb{R}$ is $(C,\alpha)$-Hölder if
$|\phi(x)-\phi(y)|\leq Cd(x,y)^{\alpha}.$
We find the following bound for the internal radius:
Theorem $1.$
Let $f\in\operatorname{Diff}^{1+}$ be a diffeomorphism admitting a dominated splitting. Let $\mu$ be an invariant measure. Then, $\mu$-almost every $x$,
(1.2) $\liminf_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}R_{\pm}(f^{k}(x))\geq\left(\frac{|\operatorname{\text{LE}}_{\pm}(x)|}{C}\right)^{\frac{1}{\alpha}}$
where $\|Df_{-}(x)\|$ and $m(Df_{+}(x))$ are $(C,\alpha)$-Hölder.
The variation of $x\mapsto E^{\pm}_{x}$ is Hölder, and this cannot be improved by increasing the differentiability of $f$.
Let $p$ be a hyperbolic periodic point. The ergodic homoclinic class $\operatorname{Phc}(p)$ of $p$ is defined as the intersection of the following two invariant sets:
$\operatorname{Phc}^{-}(p)=\{x:W^{-}(x)\pitchfork W^{+}(o(p))\neq\emptyset\},\quad\text{and}$
$\operatorname{Phc}^{+}(p)=\{x:W^{+}(x)\pitchfork W^{-}(o(p))\neq\emptyset\},$
where $o(p)$ denotes the orbit of pp.
These sets were introduced in [HHTU11].
2. Some notation and background
Theorem 2.1 (Pesin Stable Manifold Theorem [Pes76]).
Let $f\in\operatorname{Diff}^{1+}(M)$. Let $\mu$ be an invariant measure such that $\operatorname{\text{LE}}_{+}(\mu)>0$. Then for each sufficiently small $r>0$, there exists a measurable set $A_{r}$ with $\mu(A_{r})>0$ such that:
$R_{+}(x)\geq r\qquad\forall x\in A_{r}.$
An analogous statement holds for $R_{-}(x)$ if $\operatorname{\text{LE}}_{-}(\mu)<0$
We will state this classic lemma for later use:
Lemma 2.2 (Kac’s lemma).
Let $f\in\operatorname{Diff}(M)$, $\mu$ an ergodic invariant probability measure, $\psi\in L^{1}(\mu)$, AA a measurable set with $\mu(A)>0$. Define
(2.3) $\phi_{A}(x)=\min\{n>1:f^{n}(x)\in A\}.$
Then
$\int\psi d\mu=\int_{A}\sum_{k=0}^{\phi_{A}(x)-1}\psi(f^{k}(x))d\mu.$
Proof.
An interesting reference for this lemma is the unpublished notes [Sar23, Theorem 1.7]. ∎
The following criterion was introduced in [HHTU11]:
Theorem 2.3 (Criterion for ergodicity [HHTU11]).
Let $f\in\operatorname{Diff}^{1+}$ be a volume-preserving diffeomorphism and $p$ be a hyperbolic periodic point for $f$. If $m(\operatorname{Phc}^{-}(p))>0$ and $m(\operatorname{Phc}^{+}(p))>0$, then
$\operatorname{Phc}^{+}(p)\stackrel{{\scriptstyle\rm{o}}}{{=}}\operatorname{Phc}^{-}(p)\stackrel{{\scriptstyle\rm{o}}}{{=}}\operatorname{Phc}(p)$
is a hyperbolic ergodic component of $m$.
As a consequence, the Katok’s closing lemma [Kat80] allows us to write the results in the well-known Pesin work [Pes77] as:
Theorem 2.4 (Pesin’s spectral decomposition theorem).
Let $f\in\operatorname{Diff}^{1+}(M)$ be a volume-preserving diffeomorphism. Let $\operatorname{Nuh}(f)$ be the set of points without zero Lyapunov exponents. Then there exists a sequence of hyperbolic periodic points $p_{n}$ such that
$\operatorname{Nuh}(f)\stackrel{{\scriptstyle\rm{o}}}{{=}}\bigcup_{n\in\mathbb{N}}\operatorname{Phc}(p_{n}),$
where $\operatorname{Phc}(p_{n})$ are hyperbolic ergodic components of the volume measure.
If we call
$\Gamma^{\pm}(p)=\{x:W^{\pm}(x)\pitchfork W^{\mp}(p)\neq\emptyset\}$
and $\Gamma(p)=\Gamma^{+}(p)\cap\Gamma^{-}(p)$, then $f(\Gamma(f^{k}(p_{n})))=\Gamma(f^{k+1}(p_{n})$), and
$f^{\operatorname{per}(p)}|_{\Gamma(p)}\quad\text{is Bernoulli}.$
3. Average internal radius in terms of the Lyapunov exponents
For each $r>0$, define the sets
$G^{\pm}(r)=\{x:R_{\pm}(x)\geq r\}.$
Theorem 1.1 follows immediately from the following theorem.
Theorem 3.1.
Let $f\in\operatorname{Diff}^{1+}(M)$ have a dominated splitting and let $\mu$ be an ergodic invariant probability measure such that $\operatorname{\text{LE}}_{+}(\mu)>0$. Then $\mu$-almost every $x\in M$, for any choice of initial internal radius $r_{0}(x)>0$ such that $x\in G^{+}(r_{0}(x))$, there is a sequence $r_{k}(x)$ (defined in (3.4)) that satisfies:
  1. (1)
    $f^{k}(x)\in G^{+}(r_{k}(x))$ for all $k\in\mathbb{N}$,
  2. (2)
    $W^{+}_{r_{k}(x)}(f^{k}(x))\subset f(W^{+}_{r_{k-1}(x)}(f^{k-1}(x)))$ for all $k\in\mathbb{N}$,
  3. (3)
    $\operatorname{\text{LE}}_{+}(\mu)\leq C\liminf_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}r_{k}(x)^{\alpha},$
where $C,\alpha>0$ are constants such that $\log\psi_{+}$ is $(C,\alpha)$-Hölder.
We denote by $B_{\varepsilon}(x)$ the closed Riemannian ball of radius $\varepsilon>0$ centered at $x$.
Definition 3.2.
For all $x\in M$, $N\in\mathbb{N}_{0}$, and $\varepsilon>0$, define:
  1. (1)
    $\psi^{N}_{+}(x)=m(Df^{N}_{+}(x))$,
  2. (2)
    $\psi^{N}_{-}(x)=\|Df^{N}_{-}(x)\|$.
  3. (3)
    $\psi^{N}_{+}(x,\varepsilon)=\min\{\psi^{N}_{+}(y):y\in B_{\varepsilon}(x)\}$,
    $\psi^{N}_{-}(x,\varepsilon)=\max\{\psi^{N}_{-}(y):y\in B_{\varepsilon}(x)\}.$
When $N=1$, we omit 1 from the notation.
Before getting into the proof, let us see the following lemma:
Lemma $3.$
Let $f\in\operatorname{Diff}^{1}(M)$ with a dominated splitting. Let $r>0$ be such that $x\in G^{+}(r)$. If $m_{0}=\psi_{+}(x,r)$, then $f(x)\in G^{+}(m_{0}r)$. Moreover, $W_{m_{0}r}^{+}(f(x))\subset f(W^{+}_{r}(x))$.
Proof.
It is enough to see that $W^{+}_{m_{0}r}(f(x))\subset f(W^{+}_{r}(x))$. Consider a smooth path $\alpha:[0,1]\to W^{+}_{r}(x)$ such that $\alpha(0)=x$ and $\alpha(1)\in\partial W^{+}_{r}(x)$. Then
$\operatorname{length}(f\circ\alpha)=\int_{0}^{1}\|Df(\alpha(t))\alpha^{\prime}(t)\|dt\geq m_{0}\operatorname{length}(\alpha)\geq m_{0}r.$
Proof of Theorem 3.1.
Let $x\in M$ be any point such that $R_{+}(x)>0$. Choose any $r_{0}>0$ such that $r_{0}\leq R_{+}(x)$, and for each $k\in\mathbb{N}$ define inductively:
(3.4) $r_{k}=\psi_{+}(f^{k-1}(x),r_{k-1})r_{k-1}=m_{k-1}r_{k-1}$
By Lemma 3.3 $f^{k}(x)\in G^{+}(r_{k})$ for all $k\in\mathbb{N}_{0}$. Since $\log\psi_{+}$ is $(C,\alpha)$-Hölder, we have for all $N\in\mathbb{N}$:
(3.5) $\frac{1}{N}\log\frac{r_{N}}{r_{0}}=\frac{1}{N}\sum_{k=0}^{N-1}\log\psi_{+}(f^{k}(x),r_{k})\geq\frac{1}{N}\sum_{k=0}^{N-1}\log\psi_{+}(f^{k}(x))-\frac{C}{N}\sum_{k=0}^{N-1}r_{k}^{\alpha}.$
Claim 3.4.
For each $\delta>0$ and $\mu$-almost every $x\in M$ there exists $N(x,\delta)\in\mathbb{N}$ such that for all $n\geq N(x,\delta)$
$\operatorname{\text{LE}}_{+}(\mu)\leq\delta+\frac{C}{n}\sum_{k=0}^{n-1}r_{k}^{\alpha}$
Proof of Claim 3.4.
Let $L=\max\{\log\psi_{+}(x):x\in M\}>0$. Consider $N(x,\delta)>0$ such that for all $n\geq N(x,\delta)$
$\frac{1}{n}\sum_{k=0}^{n-1}\log\psi_{+}(f^{k}(x))>\operatorname{\text{LE}}_{+}(\mu)-\frac{\delta}{2},\quad\text{and}$
$\operatorname{\text{LE}}_{+}(\mu)\leq\frac{C}{n}\left(r_{0}e^{\frac{\delta}{2}n-L}\right)^{\alpha}.$
If for some $n\geq N(x,\delta)$ we had
(3.6) $\operatorname{\text{LE}}_{+}(\mu)>\delta+\frac{C}{n}\sum_{k=0}^{n-1}r_{k}^{\alpha},$
then by our choice of $N(x,\delta)$, we would have
$\frac{1}{n}\sum_{k=0}^{n-1}\log\psi_{+}(f^{k}(x))-\frac{C}{n}\sum_{k=0}^{n-1}r_{k}^{\alpha}>\frac{\delta}{2}.$
Inequality (3.5) then would imply
$r_{n}\geq e^{\frac{\delta}{2}n}r_{0}.$
Now, from Formula (3.4) we would have
$r_{n-k}=\frac{r_{n}}{m_{n-1}\cdots m_{n-k}}\geq r_{0}e^{\frac{\delta}{2}n-kL}.$
Then
$\sum_{k=0}^{n-1}r_{k}^{\alpha}\geq(r_{0}e^{\frac{\delta}{2}n})^{\alpha}\sum_{k=1}^{n}e^{-\alpha kL}.$
Our assumption (3.6) then would yield
$\operatorname{\text{LE}}_{+}(\mu)>\delta+\frac{C}{n}\left(r_{0}e^{\frac{\delta}{2}n-L}\right)^{\alpha},$
contradicting our choice of $N(x,\delta)$. This proves the claim. ∎
The claim implies item (3) and Theorem 3.1.∎
Remark 3.5.
If $R_{+}(x)=\infty$ for a measurable positive measure set $A\subset M$, it follows from Lemma 3.3 that $R_{+}(x)=\infty$ on $f(A)$. Since $\mu$ is ergodic, $R_{+}(x)=\infty$ for $\mu$-almost every $x$.
Theorem 3.6.
Let $f\in\operatorname{Diff}^{1+}(M)$ be a diffeomorphism admitting a dominated splitting. Let $\mu$ be an ergodic invariant probability measure such that $\operatorname{\text{LE}}_{+}(\mu)>0$. Then, there exists an $L^{1}(\mu)$ function $r_{+}:M\to(0,\infty)$ such that $x\in G^{+}(r_{+}(x))$ for $\mu$-almost every $x$, and
(3.7) $\operatorname{\text{LE}}_{+}(\mu)\leq\int\log\frac{r_{+}(x)}{r_{0}}d\mu+C\int r_{+}(x)^{\alpha}d\mu$
where $C,\alpha>0$ are constants such that $\log\psi_{+}$ is $(C,\alpha)$ -Hölder.
Proof.
Let $A_{r_{0}}$ be such that $\mu(A_{r_{0}})>0$, where $A_{r_{0}}$ is as in Theorem 2.1. For almost every $x\in M$, call $\phi:=\phi_{A_{r_{0}}}$ the measurable return function defined in (2.3) for the set $A_{r_{0}}$ and for $f^{-1}$ (do not confuse with $n(x)$). That is,
$\phi(x)=\min\{n>1:f^{-n}(x)\in A_{r_{0}}\}.$
For all $x\in A_{r_{0}}$, define $r_{+}(x):=r_{0}>0$, an internal radius of $W^{+}(x)$. For all $x\in\phi^{-1}(\mathbb{N})$, define $r_{+}(x):=r_{\phi(x)}$, where $r_{\phi(x)}$ is the one obtained in the recursive formula (3.4), that is:
$r_{+}(x)=\prod_{k=0}^{\phi(x)-1}\psi_{+}(f^{-k}(x),r_{+}(f^{-k}(x)))r_{0}.$
It is easy to check that $r_{+}(x)$ is a measurable function and $x\in G^{+}(r_{+}(x))$ for $\mu$-almost every $x$. If $r_{+}$ is not in $L^{1}(\mu)$, one can easily take a truncation of $r_{+}$ that is in $L^{1}(\mu)$ and satisfies (3.7) and $x\in G^{+}(r_{+}(x))$. So, assume $r_{+}$ is in $L^{1}(\mu)$. Hence, by Jensen’s inequality, $r_{+}$ is in $L^{\alpha}(\mu)$.
Now, by Kac’s Lemma (Lemma 2.2), we have:
${\operatorname{\text{LE}}_{+}(\mu)-C\int r_{+}(x)^{\alpha}d\mu=}$
$=$ $\int_{\operatorname{Pb}}\sum_{k=0}^{\phi(x)-1}[\log\psi_{+}(f^{-k}(x))-Cr_{+}(f^{-k}(x))^{\alpha}]d\mu$
$\leq$ $\int_{\operatorname{Pb}}\sum_{k=0}^{\phi(x)-1}\log\psi_{+}(f^{-k}(x),r_{+}(f^{-k}(x))d\mu$
$=$ $\int_{\phi^{-1}(\mathbb{N})}\log\frac{r_{+}(x)}{r_{0}}d\mu=\int\log\frac{r_{+}(x)}{r_{0}}d\mu$
Remark 3.7.
As a corollary of Jensen’s inequality, under the assumptions of Theorem 3.6, we get
$\operatorname{\text{LE}}_{+}(\mu)\leq\log\int\frac{r_{+}(x)}{r_{0}}d\mu+C\left(\int{r_{+}(x)}d\mu\right)^{\alpha}.$
Also, if we choose
$0<r_{0}\leq\left(\frac{\operatorname{\text{LE}}_{+}(\mu)}{C}\right)^{\frac{1}{\alpha}},$
then it follows that $\int r_{+}(x)d\mu\geq r_{0}$, otherwise, we would get a contradiction with the inequality above.
Corollary 3.8.
Let $\mu$ be an ergodic measure such that $\operatorname{\text{LE}}_{+}(\mu)>0$. Then $R_{+}$ is a measurable function.
Proof.
For $\mu$-almost every $x$ and each $k\in\mathbb{Z}$ there exists $r_{+,k}\in L^{1}$ such that $r_{+,k}(f^{k}(x))=R_{+}(f^{k}(x))$, and $y\in G^{+}(r_{+,k}(y))$ $\mu$-almost every $y$. Take $r=\sup_{k\in\mathbb{Z}}r_{+,k}$. Then $r$ is a measurable function and $r(y)=R_{+}(y)$ $\mu$-almost every $y$. ∎
4. Internal radii for periodic points
The following corollary follows immediately from Theorem 3.1:
Corollary 4.1.
Under the hypothesis of Theorem 3.1, if $p$ is a periodic point such that $\operatorname{\text{LE}}_{+}(p)>0$, then
$\operatorname{\text{LE}}_{+}(p)\leq\frac{C}{\operatorname{per}(p)}\sum_{k=0}^{\operatorname{per}(p)-1}R_{+}(f^{k}(p))^{\alpha}\leq C\left(\frac{1}{\operatorname{per}(p)}\sum_{k=0}^{\operatorname{per}(p)-1}R_{+}(f^{k}(p))\right)^{\alpha}.$
Proof.
Let $r_{0}=R_{+}(p)$ and let us do the inductive procedure in (3.4). Then we obtain
$\frac{1}{N}\log\frac{r_{N}}{R_{+}(p)}=\frac{1}{N}\sum_{k=0}^{N-1}\log m_{k}\leq 0,$
where $N=\operatorname{per}(p)$. Otherwise we would obtain that $r_{N}>R_{+}(p)$, which is a contradiction. We also have $\psi_{+}(f^{k}(p),R_{+}(f^{k}(p)))\leq m_{k}$ for all $k\in[0,\operatorname{per}(p)-1]$.
Due to the $(C,\alpha)$-Hölderness of the function $\log\psi_{+}$, the following holds
$\log\psi_{+}(f^{k}(p),R_{+}(f^{k}(p)))\geq\log\psi_{+}(f^{k}(p))-CR_{+}(f^{k}(p))^{\alpha}.$
The result then follows. ∎
Proposition 4.2.
Let $p$ be a hyperbolic periodic point of period NN, then
$R_{+}(p)\geq d(p,M\setminus A^{+}(N)),$
where $A^{+}(N)=\{x\in M:\log\psi^{N}_{+}(x)>0\}$.
Proof.
It follows from Lemma 3.3 for $f^{N}$ that $\psi^{N}_{+}(p,R_{+}(p))\leq 1$, for otherwise we would obtain a contradiction with our choice of $R_{+}(p)$. This implies that $W^{+}_{R_{+}(p)}(p)\cap(M\setminus A^{+}(N))\neq\emptyset$. This implies that $d(p,M\setminus A^{+}(N))\leq R_{+}(p)$.
Remark 4.3.
  1. (1)
    Maybe it is handy to note the following: if $\overline{W^{+}_{\varepsilon}(p)}\subset A^{+}(\operatorname{per}(p))$, then $R_{+}(p)>\varepsilon$.
  2. (2)
5. Time bounds
Definition 5.1 (Pesin blocks).
Given $f\in\operatorname{Diff}^{1}(M)$ admitting a $\gamma$-dominated splitting, the Pesin blocks are the sets of the form:
$\operatorname{Pb}^{+}_{N}(\gamma)=\left\{x\in M^{+}:\frac{1}{n}\sum_{k=0}^{n-1}\log\psi_{+}f^{k}(x))\geq\gamma/2\quad\forall n\geq N\right\},$
$\operatorname{Pb}^{-}_{N}(\gamma)=\left\{x\in M^{-}:\frac{1}{n}\sum_{k=0}^{n-1}\log\psi_{-}(f^{k}(x))\leq-\gamma/2\quad\forall n\geq N\right\}.$
Pesin blocks are closed sets where there is “uniform hyperbolicity”, but at the cost of not being invariant.
Proposition 5.2.
For all $x\in\operatorname{Pb}^{+}_{N}(\gamma)\cap G^{+}(R_{0}e^{-K\gamma/4})$,
$\sup_{0\leq k\leq\max(K,N)}R_{+}(f^{k}(x))\geq R_{0},$
where $R_{0}=(\frac{\gamma}{4C})^{\frac{1}{\alpha}}$.
Proof.
Suppose
$\sup_{0\leq k\leq\max(K,N)}R_{+}(f^{k}(x))<R_{0}$
and call $T=\max(K,N)$. Then
$\frac{1}{T}\log\frac{R_{+}(f^{T}(x))}{R^{+}(x)}\geq\frac{\gamma}{2}-\frac{C}{T}\sum_{k=0}^{T-1}\sup R_{+}(f^{k}(x))^{\alpha}\geq\frac{\gamma}{4}.$
This implies that
$R_{+}(f^{T}(x))\geq R^{+}(x)\exp\left(\frac{\gamma}{4}K\right)\geq R_{0}.$
This produces a contradiction. ∎